# On momoid graded semihereditary rings

### v1

@@ Abstract.
In order to study graded left hereditary and left semihereditary rings graded by a cancelation monoid in terms of their modules, we need to revisit graded free, projective, injective, and flat modules and provide graded versions of specific results concerning these modules, like Baer’s criterion on injectivity and Lazard’s theorem on flatness. Then, among other things, we can give some characterization of graded left hereditary and left semihereditary rings, in particular, of graded-Prüfer and graded-Dedekind domains.
@@ 1. Introduction
After Kaplansky’s work on Dedekind and Prüfer domains [11], Cartan and Eilenberg defined the notion of hereditary and semihereditary rings in [7]. Let $R$ be an associative ring with nonzero identity. Recall that $R$ is left hereditary if every left ideal is a projective $R$-module. It is known that $R$ is hereditary if and only if each submodule of a projective left $R$-module is projective if and only if each quotient of an injective left $R$-module is injective [7, Page 14]. A hereditary domain is a Dedekind domain. It was proved that $R$ is a Dedekind domain if and only if every divisible $R$-module is injective [18, Theorem 4.24].
@@
Recall also that $R$ is left semihereditary if every finitely generated left ideal is a projective $R$-module. It is known that $R$ is left semihereditary if and only if $R$ is left coherent, and each submodule of a flat left $R$-module is flat if and only if each torsion-free right $R$-module is flat [18, Theorem 4.32]. A semihereditary domain is a Prüfer domain. It is shown that $R$ is a Prüfer domain if and only if every finitely generated torsion-free $R$-module is projective [18, Theorem 4.34].
@@
The graded analogue of Dedekind and Prüfer domains was defined and studied by several authors (see [14, 2, 4, 19]). This paper aims to define and study the graded analogues of hereditary and semihereditary rings. Let $\Gamma$ be a cancelation monoid, and $R=\bigoplus_{\alpha\in\Gamma}R_{\alpha}$ be a $\Gamma$-graded ring. To this purpose, one needs to study $\Gamma$-graded free, projective, injective, and flat modules. These modules were extensively studied in case where $\Gamma=\mathbb{Z}$ in [9, 6], and in case where $\Gamma$ is a group in [15]. The first two of them were investigated in case where $\Gamma$ is a cancelation monoid in [13]. Then, we examine $\Gamma$-graded injective and flat modules where $\Gamma$ is a cancelation monoid.
@@
The organization of this paper is as follows. In the remainder of the introduction and Section 2, we recall some preliminaries on graded rings and modules. In Section 3, after reviewing some results on the graded free and projective modules, we investigate the graded divisible, injective, and flat modules. We provide a graded analogue of Baer’s Theorem and show that every graded injective module is graded divisible and every $R$-module admits a graded injective envelope. Then, we define and study graded flat modules and extend Lazard’s Theorem to the monoid-graded case. More precisely, it is shown that any graded flat module is a direct limit of finitely generated graded free modules. In section 4, we study graded hereditary rings. A $\Gamma$-graded ring $R=\bigoplus_{\alpha\in\Gamma}R_{\alpha}$ is called graded left hereditary if every left homogeneous ideal is projective. Then, among other things, we prove the graded version of the Cartan-Eilenberg Theorem: $R$ is graded left hereditary if and only if every graded submodule of a graded projective $R$-module is graded projective if and only if every quotient of a graded injective $R$-module is graded injective. Then we show that a graded (commutative) domain is graded-Dedekind if and only if every graded-divisible left $R$-module is graded-injective. Finally, in Section 5, we study graded semihereditary rings. A $\Gamma$-graded ring $R$ is called graded left semihereditary if every finitely generated left homogeneous ideal is projective. Among other things, it is shown that $R$ is graded left semihereditary if and only if every finitely generated graded submodule of a graded projective left $R$-module is graded projective if and only if $R$ is graded left coherent and every homogeneous left ideal is graded flat. It is also shown that a graded integral domain $R$ is a graded-Prüfer domain if and only if every finitely generated torsion-free graded $R$-module is graded projective.
@@ 1.1. Graded Rings and modules
Let $R=\bigoplus_{\alpha\in\Gamma}R_{\alpha}$ be a $\Gamma$-graded ring. By a $\Gamma$-graded $R$-module $M$, we mean a left $R$-module which has a direct sum decomposition $M=\bigoplus_{\alpha\in\Gamma}M_{\alpha}$, where each $M_{\alpha}$ is an additive subgroup of $M$, such that $R_{\alpha}M_{\beta}\subseteq M_{\alpha\beta}$ for all $\alpha,\beta\in\Gamma$. Similarly, one can define a $\Gamma$-graded right $R$-module. For each $\alpha\in\Gamma$, we call $0\neq x\in M_{\alpha}$ a homogeneous element of degree $\alpha$ and write $\deg(x)=\alpha$. By [13, Lemma 1.1], if $\{x_{i}\}_{i\in J}$ is a homogeneous generating set of $M$ with $\deg(x_{i})=\alpha_{i}$, $i\in J$, then
@@
Let $M=\bigoplus_{\alpha\in\Gamma}M_{\alpha}$ and $N=\bigoplus_{\alpha\in\Gamma}N_{\alpha}$ be $\Gamma$-graded left $R$-modules. If $\varphi:M\to N$ is an $R$-module homomorphism such that $\varphi(M_{\alpha})\subseteq N_{\alpha}$ for all $\alpha\in\Gamma$, then we call $\varphi$ a homogeneous homomorphism. In the case that $\varphi:M\to N$ is a homogeneous homomorphism, it is easy to see that $\ker(\varphi)$ and $\operatorname{im}(\varphi)$ are graded submodules of $M$ and $N$, respectively.
@@ 2. Preliminaries
Again, by using the cancelation law on $\Gamma$, one notices that the functors $\sideset{{}^{*}}{{}_{R}}{\operatorname{Hom}}(-,N)$ and $\sideset{{}^{*}}{{}_{R}}{\operatorname{Hom}}(M,-)$ are left exact on the category of $\Gamma$-graded left $R$-modules. In general, $\sideset{{}^{*}}{{}_{R}}{\operatorname{Hom}}(M,N)$ may not equal to $\operatorname{Hom}_{R}(M,N)$, but under certain conditions we have $\sideset{{}^{*}}{{}_{R}}{\operatorname{Hom}}(M,N)=\operatorname{Hom}_{R}(M,N)$, for instance, for the case that $\Gamma=G$ is a group, if $M$ is finitely generated ([15, Corollary 2.4.4]), or if both $M$ and $N$ have finite support ([15, Corollary 2.4.5]), in particular in the case where the grading group $G$ is finite.
@@ 3.1. Graded-projective modules
Recall that a $\Gamma$-graded left $R$-module $F=\bigoplus_{\alpha\in\Gamma}F_{\alpha}$ is called graded free (for short, gr-free) if $F$ is a free $R$-module (in the usual sense) but with a homogeneous free $R$-basis; namely, $F$ has a free $R$-basis $\{e_{i}\}_{i\in J}$ consisting of homogeneous elements. Given any free $R$-module $F=\bigoplus_{i\in J}Re_{i}$ and an arbitrarily chosen family $\{\alpha_{i}\}_{i\in J}$ of elements of $\Gamma$, one can give a $\Gamma$-graduation on $F$ such that $\{e_{i}\}_{i\in J}$ will be a homogeneous basis with $\deg(e_{i})=\alpha_{i}$. This shows that if $M$ is a $\Gamma$-graded left $R$-module with homogeneous generating set $\{x_{i}\}_{i\in J}$, and the gr-free $R$-module $F=\bigoplus_{i\in J}Re_{i}$, with $\deg(e_{i})=\deg(x_{i})$, then the map $\varphi:F\to M$ defined by $\varphi(e_{i})=x_{i}$ is a homogeneous epimorphism. Hence, every nonzero $\Gamma$-graded left $R$-module is a graded homomorphic image of some gr-free $R$-module (see [13, Page 2698] for details).
@@
$(2)\Rightarrow(3)$ As we have already seen $\sideset{{}^{*}}{{}_{R}}{\operatorname{Hom}}(P,-)$ is left exact. The right exactness follows immediately from the definition of gr-projective modules.
@@
Assume that $M$ is a finitely presented $\Gamma$-graded left $R$-module. Let $F_{1}\to F_{0}\stackrel{{\scriptstyle f}}{{\to}}M\to 0$ be a finite presentation of $M$ (in the usual sense); so that $0\to\ker(f)\to F_{0}\stackrel{{\scriptstyle f}}{{\to}}M\to 0$ is exact. Note that $M$ is finitely generated; so we can choose a finite set of generators $x_{1},\ldots,x_{n}$, such that each $x_{i}$ is a homogeneous element of degree $\alpha_{i}$. Let $F^{\prime}_{0}:=\oplus_{i=1}^{n}Re_{i}$ be the gr-free $\Gamma$-graded left $R$-module with homogeneous basis $\{e_{1},\ldots,e_{n}\}$, $\deg(e_{i})=\alpha_{i}$. Define the homogeneous homomorphism $f^{\prime}:F^{\prime}_{0}\to M$ by $e_{i}\mapsto x_{i}$. Hence, there is an exact sequence $0\to\ker(f^{\prime})\to F^{\prime}_{0}\stackrel{{\scriptstyle f^{\prime}}}{{\to}}M\to 0$. It follows from Schanuel’s Lemma [18, Proposition 3.12] that $\ker(f)\oplus F^{\prime}_{0}\cong\ker(f^{\prime})\oplus F_{0}$. Since $\ker(f)$ is finitely generated, we see that $\ker(f^{\prime})$ is finitely generated, too. Hence $\ker(f^{\prime})$ is an image of a finitely generated gr-free left $R$-module $F^{\prime}_{1}$. Therefore, $M$ is finitely presented in the category of $\Gamma$-graded left $R$-modules.
@@ 3.2. Graded-injective modules
The notion of graded injective modules was investigated in [9, 6] in the case where $\Gamma=\mathbb{Z}$ and in [15] in the case where $\Gamma=G$ is a group. It is known that the notions of graded injective and (usual) injective are not the same. For example, over an arbitrary field $k$, consider the Laurent polynomial ring $R=k[X,X^{-1}]$ with $\mathbb{Z}$-gradation. Then by [15, Remark 2.3.3], the left $R$-module $R$ is graded injective but not injective.
A $\Gamma$-graded left $R$-module $E$ is said to be a graded injective module (for short, gr-injective module) if for every homogeneous homomorphisms $f:M\to E$ and $g:M\to N$ with $g$ is a homogeneous injection, there exists a homogeneous homomorphism $h:N\to E$ such that the diagram
@@
We now provide some useful characterization of gr-injective modules, which will be used in the sequel (see [15, Corollary 2.4.8] when $\Gamma=G$ is a group).
@@
Assume that $M$ is gr-injective, $I$ is a homogeneous left ideal of $R$, and $f:I\to M$ is a homogeneous homomorphism of degree $\alpha\in\Gamma$. There is a gr-free left $R$-module $F=Re$ with homogeneous basis $\{e\}$ of degree $\alpha$. Consider the following diagram
@@
commutes. Partially order $\Sigma$ by $(B^{\prime\prime},g^{\prime\prime})\leq(B^{\prime},g^{\prime})$ when $B^{\prime\prime}\subseteq B^{\prime}$ and $g^{\prime}\mid_{B^{\prime\prime}}=g^{\prime\prime}$. Then $\Sigma$ is nonempty and inductive. Let $(B^{\prime},g^{\prime})$ be a maximal element of $\Sigma$. We claim that $B^{\prime}=B$. Otherwise, choose a homogeneous element $x\in B\setminus B^{\prime}$ of degree $\alpha$. Let $I=\{r\in R\mid rx\in B^{\prime}\}$ which is a homogeneous left ideal of $R$. Define $\overline{f}:I\to M$ by $\overline{f}(r)=g^{\prime}(rx)$ for $r\in I$. It is clear that $\overline{f}$ is a homogeneous homomorphism of degree $\alpha$. By assumption, there is $m_{\alpha}\in M_{\alpha}$ such that $\overline{f}(r)=rm_{\alpha}$ for $r\in I$. Set $B^{\prime\prime}:=B^{\prime}+Rx$ and define $g^{\prime\prime}:B^{\prime\prime}\to M$ by $g^{\prime\prime}(b+rx)=g^{\prime}(b)+rm_{\alpha}$. It can be see that $g^{\prime\prime}$ is a well defined homomorphism. We show that $g^{\prime\prime}$ is homogeneous. Indeed, assume that $\gamma\in\Gamma$ and $b^{\prime\prime}\in B^{\prime\prime}_{\gamma}$. Then there are $b\in B^{\prime}_{\gamma}$ and $r\in R_{\beta}$ with $\beta\alpha=\gamma$ such that $b^{\prime\prime}=b+rx$. Hence $rm_{\alpha}$ is homogeneous of degree $\beta\alpha=\gamma$; so $g^{\prime\prime}(b^{\prime\prime})=g^{\prime}(b)+rm_{\alpha}$ is homogeneous of degree $\gamma$. This means that $(B^{\prime\prime},g^{\prime\prime})$ properly extends $(B^{\prime},g^{\prime})$, contradicting the maximality of $(B^{\prime},g^{\prime})$. ∎
@@
    Every short exact sequence of $\Gamma$-graded left $R$-module and homogeneous homomorphisms $0\to E\to M\to N\to 0$ splits via a homogeneous homomorphism.
@@
It is well known that every (not necessarily graded) injective $R$-module is divisible, and the converse is true if $R$ is a Dedekind domain [18, Theorem 4.24]. The notion of a graded divisible module was introduced in [10, Page 47] for group-graded modules. This notion can be generalized to monoid-graded modules as in the following. Recall that a nonzero element $a\in R$ is a right zero divisor if there exists a nonzero element $b\in R$ such that $ba=0$. The element $a\in R$ is called to be a right nonzero divisor if it is not a right zero divisor.
@@
It is clear that every graded and divisible $R$-module is gr-divisible. But the converse is no longer true. For, over an arbitrary field $k$, consider the Laurent polynomial ring $R=k[X,X^{-1}]$ with $\mathbb{Z}$-gradation. Then by [15, Remark 2.3.3], the left $R$-module $R$ is gr-injective but not injective. Hence, it is gr-divisible by the following proposition but not divisible since $R$ is a Dedekind domain.
@@
Let $M$ be a $\Gamma$-graded left $R$-module, and let $N$ be a graded submodule of $M$. The extension $N\subseteq M$ is called a graded-essential extension (for short, gr-essential extension) if $U\cap N\neq 0$ for any graded submodule $0\neq U\subseteq M$. If in addition $N\neq M$, the extension is called a proper gr-essential extension. It can be seen as in the case where $\Gamma=\mathbb{Z}$ (see [9, Lemma 1.1]) or $\Gamma=G$ a group (see [15, Proposition 2.3.5]), one can show that $N\subseteq M$ is a gr-essential extension if and only if it is an essential extension.
@@
Conversely, suppose that $N$ has no proper gr-essential extension. Given a homogeneous monomorphism $\varphi:U\to V$ and a homogeneous homomorphism $f:U\to M$, we want to construct a homogeneous homomorphism $h:V\to N$ such that $f=h\phi$. We consider the pushout diagram
@@
Any $\Gamma$-graded left $R$-module $M$ admits a gr-injective envelope, and any two gr-injective envelopes are isomorphic as $\Gamma$-graded left $R$-module.

### v3

# On monoid graded semihereditary rings

@@ Abstract.
In order to study graded left hereditary and left semihereditary rings graded by a cancelation monoid in terms of their modules, we need to revisit graded free, projective, injective, and flat modules and provide graded versions of specific results concerning these modules, like Baer’s criterion on injectivity and Lazard’s theorem on flatness. Then, among other things, we can give some characterizations of graded left hereditary and left semihereditary rings, in particular, of graded-Prüfer and graded-Dedekind domains.
@@ 1. Introduction
After Kaplansky’s work on Dedekind and Prüfer domains [11], Cartan and Eilenberg defined the notion of hereditary and semihereditary rings in [7]. Let $R$ be an associative ring with nonzero identity. Recall that $R$ is left hereditary if every left ideal is a projective $R$-module. It is known that $R$ is left hereditary if and only if each submodule of a projective left $R$-module is projective if and only if each quotient of an injective left $R$-module is injective [7, Page 14]. A commutative hereditary integral domain is a Dedekind domain. It was proved that $R$ is a Dedekind domain if and only if every divisible $R$-module is injective [18, Theorem 4.24].
@@
Recall also that $R$ is left semihereditary if every finitely generated left ideal is a projective $R$-module. It is known that $R$ is left semihereditary if and only if $R$ is left coherent, and each submodule of a flat left $R$-module is flat if and only if each torsion-free right $R$-module is flat [18, Theorem 4.32]. A commutative semihereditary integral domain is a Prüfer domain. It is shown that $R$ is a Prüfer domain if and only if every finitely generated torsion-free $R$-module is projective [18, Theorem 4.34].
@@
The graded analogue of Dedekind and Prüfer domains was defined and studied by several authors (see [14, 2, 4, 19]). This paper aims to define and study the graded analogues of left hereditary and semihereditary rings. Let $\Gamma$ be a cancelation monoid, and $R=\bigoplus_{\alpha\in\Gamma}R_{\alpha}$ be a $\Gamma$-graded ring. To this purpose, one needs to study $\Gamma$-graded free, projective, injective, and flat modules. These modules were extensively studied in case where $\Gamma=\mathbb{Z}$ in [9, 6], and in case where $\Gamma$ is a group in [15]. The first two of them were investigated in case where $\Gamma$ is a cancelation monoid in [13]. Then, in this paper, we examine $\Gamma$-graded injective and flat modules where $\Gamma$ is a cancelation monoid.
@@
The organization of this paper is as follows. In the remainder of the introduction and Section 2, we recall some preliminaries on graded rings and modules. In Section 3, after reviewing some results on the graded free and projective modules, we investigate the graded divisible, injective, and flat modules. We provide a graded analogue of Baer’s Theorem and show that every graded injective module is graded divisible and every $R$-module admits a graded injective envelope. Then, we define and study graded flat modules and extend Lazard’s Theorem to the monoid-graded case. More precisely, it is shown that any graded flat module is a direct limit of finitely generated graded free modules. In Section 4, we study graded left hereditary rings. A $\Gamma$-graded ring $R=\bigoplus_{\alpha\in\Gamma}R_{\alpha}$ is called graded left hereditary if every left homogeneous ideal is projective. Then, among other things, we prove the graded version of the Cartan-Eilenberg Theorem: $R$ is graded left hereditary if and only if every graded submodule of a graded projective $R$-module is graded projective if and only if every quotient of a graded injective $R$-module is graded injective. Then we show that a commutative graded integral domain is graded-Dedekind if and only if every graded-divisible left $R$-module is graded-injective. Finally, in Section 5, we study graded left semihereditary rings. A $\Gamma$-graded ring $R$ is called graded left semihereditary if every finitely generated left homogeneous ideal is projective. Among other things, it is shown that $R$ is graded left semihereditary if and only if every finitely generated graded submodule of a graded projective left $R$-module is graded projective if and only if $R$ is graded left coherent and every homogeneous left ideal is graded flat. It is also shown that a commutative graded integral domain $R$ is a graded-Prüfer domain if and only if every finitely generated torsion-free graded $R$-module is graded projective.
@@ 1.1. Graded Rings and modules
Let $R=\bigoplus_{\alpha\in\Gamma}R_{\alpha}$ be a $\Gamma$-graded ring. By a $\Gamma$-graded left $R$-module $M$, we mean a left $R$-module which has a direct sum decomposition $M=\bigoplus_{\alpha\in\Gamma}M_{\alpha}$, where each $M_{\alpha}$ is an additive subgroup of $M$, such that $R_{\alpha}M_{\beta}\subseteq M_{\alpha\beta}$ for all $\alpha,\beta\in\Gamma$. Similarly, one can define a $\Gamma$-graded right $R$-module. For each $\alpha\in\Gamma$, we call $0\neq x\in M_{\alpha}$ a homogeneous element of degree $\alpha$ and write $\deg(x)=\alpha$. By [13, Lemma 1.1], if $\{x_{i}\}_{i\in J}$ is a homogeneous generating set of $M$ with $\deg(x_{i})=\alpha_{i}$, $i\in J$, then
@@
Let $M=\bigoplus_{\alpha\in\Gamma}M_{\alpha}$ and $N=\bigoplus_{\alpha\in\Gamma}N_{\alpha}$ be $\Gamma$-graded left $R$-modules. If $\varphi:M\to N$ is an $R$-module homomorphism such that $\varphi(M_{\alpha})\subseteq N_{\alpha}$ for all $\alpha\in\Gamma$, then we call $\varphi$ a homogeneous homomorphism. In the case that $\varphi:M\to N$ is a homogeneous homomorphism, it is easy to see that $\ker(\varphi)$ and $\operatorname{im}(\varphi)$ are graded left submodules of $M$ and $N$, respectively.
@@ 2. Preliminaries
Again, by using the cancelation law on $\Gamma$, one notices that the functors $\sideset{{}^{*}}{{}_{R}}{\operatorname{Hom}}(-,N)$ and $\sideset{{}^{*}}{{}_{R}}{\operatorname{Hom}}(M,-)$ are left exact on the category of $\Gamma$-graded left $R$-modules. In general, $\sideset{{}^{*}}{{}_{R}}{\operatorname{Hom}}(M,N)$ may not equal to $\operatorname{Hom}_{R}(M,N)$, but under certain conditions we have $\sideset{{}^{*}}{{}_{R}}{\operatorname{Hom}}(M,N)=\operatorname{Hom}_{R}(M,N)$, for instance, for the case that $\Gamma$ is a group, if $M$ is finitely generated ([15, Corollary 2.4.4]), or if both $M$ and $N$ have finite support ([15, Corollary 2.4.5]), in particular in the case where the grading group $\Gamma$ is finite.
@@ 3.1. Graded-projective modules
Recall that a $\Gamma$-graded left $R$-module $F$ is called graded free (for short, gr-free) if it has an $R$-basis consisting of homogeneous elements. One notices that there is a graded and free modules that is not gr-free [15, Page 21]; hence gr-free is a stronger property than “graded plus free”. Given any free $R$-module $F=\bigoplus_{i\in J}Re_{i}$ and an arbitrarily chosen family $\{\alpha_{i}\}_{i\in J}$ of elements of $\Gamma$, one can give a $\Gamma$-graduation on $F$ such that $\{e_{i}\}_{i\in J}$ will be a homogeneous basis with $\deg(e_{i})=\alpha_{i}$. This shows that if $M$ is a $\Gamma$-graded left $R$-module with homogeneous generating set $\{x_{i}\}_{i\in J}$, and the gr-free $R$-module $F=\bigoplus_{i\in J}Re_{i}$, with $\deg(e_{i})=\deg(x_{i})$, then the map $\varphi:F\to M$ defined by $\varphi(e_{i})=x_{i}$ is a homogeneous epimorphism. Hence, every nonzero $\Gamma$-graded left $R$-module is a graded homomorphic image of some gr-free $R$-module (see [13, Page 2698] for details).
@@
$(2)\Rightarrow(3)$ As we have already seen, $\sideset{{}^{*}}{{}_{R}}{\operatorname{Hom}}(P,-)$ is left exact. The right exactness follows immediately from the definition of gr-projective modules.
@@
Assume that $M$ is a $\Gamma$-graded left $R$-module which is finitely presented in the category of left $R$-modules. Let $F_{1}\to F_{0}\stackrel{{\scriptstyle f}}{{\to}}M\to 0$ be a finite presentation of $M$ (in the usual sense); so that $0\to\ker(f)\to F_{0}\stackrel{{\scriptstyle f}}{{\to}}M\to 0$ is exact. Note that $M$ is finitely generated; so we can choose a finite set of generators $x_{1},\ldots,x_{n}$, such that each $x_{i}$ is a homogeneous element of degree $\alpha_{i}$. Let $F^{\prime}_{0}:=\oplus_{i=1}^{n}Re_{i}$ be the gr-free $\Gamma$-graded left $R$-module with homogeneous basis $\{e_{1},\ldots,e_{n}\}$, $\deg(e_{i})=\alpha_{i}$. Define the homogeneous homomorphism $f^{\prime}:F^{\prime}_{0}\to M$ by $e_{i}\mapsto x_{i}$. Hence, there is an exact sequence $0\to\ker(f^{\prime})\to F^{\prime}_{0}\stackrel{{\scriptstyle f^{\prime}}}{{\to}}M\to 0$. It follows from Schanuel’s Lemma [18, Proposition 3.12] that $\ker(f)\oplus F^{\prime}_{0}\cong\ker(f^{\prime})\oplus F_{0}$. Since $\ker(f)$ is finitely generated, we see that $\ker(f^{\prime})$ is finitely generated, too. Hence $\ker(f^{\prime})$ is an image of a finitely generated gr-free left $R$-module $F^{\prime}_{1}$. Therefore, $M$ is finitely presented in the category of $\Gamma$-graded left $R$-modules.
@@ 3.2. Graded-injective modules
The notion of graded injective modules was investigated in [9, 6] in the case where $\Gamma=\mathbb{Z}$ and in [15] in the case where $\Gamma$ is a group. It is known that the notions of graded injective and (usual) injective are not the same. For example, over an arbitrary field $k$, consider the Laurent polynomial ring $R=k[X,X^{-1}]$ with $\mathbb{Z}$-gradation. Then by [15, Remark 2.3.3], the left $R$-module $R$ is graded injective but not injective.
A $\Gamma$-graded left $R$-module $E$ is said to be a graded injective module (for short, gr-injective module) if for every homogeneous homomorphisms $f:M\to E$ and $g:M\to N$ with $g$ a homogeneous injection, there exists a homogeneous homomorphism $h:N\to E$ such that the diagram
@@
We now provide some useful characterizations of gr-injective modules, which will be used in the sequel (see [15, Corollary 2.4.8] when $\Gamma$ is a group).
@@
Assume that $M$ is gr-injective, $I$ is a homogeneous left ideal of $R$, and $f:I\to M$ is a homogeneous homomorphism of degree $\alpha\in\Gamma$. There is a gr-free left $R$-module $F=Re$ with homogeneous basis $\{e\}$ of degree $\alpha$ by [13, Page 2698]. Consider the following diagram
@@
commutes. Partially order $\Sigma$ by $(B^{\prime\prime},g^{\prime\prime})\leq(B^{\prime},g^{\prime})$ when $B^{\prime\prime}\subseteq B^{\prime}$ and $g^{\prime}\mid_{B^{\prime\prime}}=g^{\prime\prime}$. Then $\Sigma$ is nonempty and inductive. Let $(B^{\prime},g^{\prime})$ be a maximal element of $\Sigma$. We claim that $B^{\prime}=B$. Otherwise, choose a homogeneous element $x\in B\setminus B^{\prime}$ of degree $\alpha$. Let $I=\{r\in R\mid rx\in B^{\prime}\}$ which is a homogeneous left ideal of $R$. Define $\overline{f}:I\to M$ by $\overline{f}(r)=g^{\prime}(rx)$ for $r\in I$. It is clear that $\overline{f}$ is a homogeneous homomorphism of degree $\alpha$. By assumption, there is $m_{\alpha}\in M_{\alpha}$ such that $\overline{f}(r)=rm_{\alpha}$ for $r\in I$. Set $B^{\prime\prime}:=B^{\prime}+Rx$ and define $g^{\prime\prime}:B^{\prime\prime}\to M$ by $g^{\prime\prime}(b+rx)=g^{\prime}(b)+rm_{\alpha}$. It can be seen that $g^{\prime\prime}$ is a well defined homomorphism. We show that $g^{\prime\prime}$ is homogeneous. Indeed, assume that $\gamma\in\Gamma$ and $b^{\prime\prime}\in B^{\prime\prime}_{\gamma}$. Then there are $b\in B^{\prime}_{\gamma}$ and $r\in R_{\beta}$ with $\beta\alpha=\gamma$ such that $b^{\prime\prime}=b+rx$. Hence $rm_{\alpha}$ is homogeneous of degree $\beta\alpha=\gamma$; so $g^{\prime\prime}(b^{\prime\prime})=g^{\prime}(b)+rm_{\alpha}$ is homogeneous of degree $\gamma$. This means that $(B^{\prime\prime},g^{\prime\prime})$ properly extends $(B^{\prime},g^{\prime})$, contradicting the maximality of $(B^{\prime},g^{\prime})$. ∎
@@
    Every short exact sequence of $\Gamma$-graded left $R$-modules and homogeneous homomorphisms $0\to E\to M\to N\to 0$ splits via a homogeneous homomorphism.
@@
It is well known that every (not necessarily graded) injective $R$-module is divisible [12, Corollary (3.17)], and a commutative integral domain $R$ is a Dedekind domain if and only if every divisible $R$-module is injective [18, Theorem 4.24]. The notion of a graded divisible module was introduced in [10, Page 47] for group-graded modules. This notion can be generalized to monoid-graded modules as in the following. Recall that a nonzero element $a\in R$ is a right zero divisor if there exists a nonzero element $b\in R$ such that $ba=0$. The element $a\in R$ is called to be a right nonzero divisor if it is not a right zero divisor.
@@
It is clear that every graded and divisible $R$-module is gr-divisible. But the converse is no longer true. For example, over an arbitrary field $k$, consider the Laurent polynomial ring $R=k[X,X^{-1}]$ with $\mathbb{Z}$-gradation. Then by [15, Remark 2.3.3], the left $R$-module $R$ is gr-injective but not injective. Hence, it is gr-divisible by the following proposition but not divisible since $R$ is a Dedekind domain (see [18, Theorem 4.24]).
@@
Let $M$ be a $\Gamma$-graded left $R$-module, and let $N$ be a graded submodule of $M$. The extension $N\subseteq M$ is called a graded-essential extension (for short, gr-essential extension) if $U\cap N\neq 0$ for any graded submodule $0\neq U\subseteq M$. If in addition $N\neq M$, the extension is called a proper gr-essential extension. As in the case where $\Gamma=\mathbb{Z}$ (see [9, Lemma 1.1]) or $\Gamma$ is a group (see [15, Proposition 2.3.5]), one can show that $N\subseteq M$ is a gr-essential extension if and only if it is an essential extension.
@@
Conversely, suppose that $N$ has no proper gr-essential extension. Given a homogeneous monomorphism $\varphi:U\to V$ and a homogeneous homomorphism $f:U\to M$, we want to construct a homogeneous homomorphism $h:V\to N$ such that $f=h\varphi$. We consider the pushout diagram
@@
Any $\Gamma$-graded left $R$-module $M$ admits a gr-injective envelope, and any two gr-injective envelopes are isomorphic as $\Gamma$-graded left $R$-modules.